OpenStudy (anonymous):

tara bought 5 1/4 yards of material to make a coustume for a school party she used 3 2/3 yards for coustume how much is left answer in simplest form tara bought 5 1/4 yards of material to make a coustume for a school party she used 3 2/3 yards for coustume how much is left answer in simplest form @Mathematics

OpenStudy (mathteacher1729):

There are a few ways of doing this problem. Which ways have you tried already?

OpenStudy (anonymous):

subtract

OpenStudy (mathteacher1729):

What problem are you running into when you try to subtract?

OpenStudy (anonymous):

5 1/4- 3 2/3

OpenStudy (mathteacher1729):

Do you know how to convert a mixed number into a fraction?

OpenStudy (anonymous):

no

OpenStudy (mathteacher1729):

Ok. That's where to begin then. :) The basic idea is " step 1 multiply the whole number (5) by the number in the denominator (4) and step 2 add that to the number in the numerator (1), then step 3 put that number over the number in the denominator (4). It's actually pretty quick. 5 + 1/4 ==> 5*4 = 20. 20 + 1 = 21 final answer = 21/4 Understanding WHY we do all that in the first place is SUPER DUPER important, otherwise there is no point in working the problem. :)

OpenStudy (anonymous):

an object that weighs one pound on earth can weigh much more than that on other planets the list below shows how more an object can weigh on other planets jupiter:2 1/2 neptune:1 1/5 what would be the weight of a 12 3/4 pound bowling ball on jupiter and neptune @Mathematics

OpenStudy (mathteacher1729):

You would take the weight of the object on earth and multiply it by the number associated with each planet. (so multiply 12 + 3/4 times 2 +1/2 to find the weight on Jupiter. FUN FACT -- Jupiter has no solid surface. It's just a gigantic ball of gas.